Solution is prepared by dissolving 50g of NaCl (MM= 58.44g/mol) in 50g of H2O (MM = 18g/mol). The volume of the solution is 63.3 ml. Find mass % of solute, m, M, N of the solution. (M-molarity, m-molality, N-normality)



1
Expert's answer
2021-04-16T05:00:37-0400

Solution:

Mass of NaCl = 50 g

The molar mass of NaCl is 58.44 g/mol

Mass of H2O = 50 g = 0.05 kg

The molar mass of H2O is 18.00 g/mol.

The volume of the solution is 63.3 ml (0.0633 L).

 

Mass % of solute (NaCl):

The mass/mass percent (%m/m) is defined as the mass of a solute divided by the mass of a solution times 100.

%m/m = (mass of solute / mass of solution) × 100%

Mass of solution = Mass of NaCl + Mass of H2O = 50 g + 50 g = 100 g

Mass of solute (NaCl) = 50 g

Hence,

Mass % of solute = (50 g / 100g) × 100% = 50%

Mass % of solute = 50%


Molarity (M):

Molarity or molar concentration is the number of moles of solute per liter of solution.

Molarity of solution = Moles of solute / Volume of solution

Moles of NaCl = Mass of NaCl / Molar mass of NaCl = 50 g / 58.44 g mol-1 = 0.8556 mol

Hence,

Molarity of solution = 0.8556 mol / 0.0633 L = 13.5166 mol/L = 13.5 M

Molarity of solution = 13.5 M

 

Molality (m):

Molality of a solution is the moles of solute divided by the kilograms of solvent.

Molality of solution = Moles of solute / Kilograms of solvent

Mass of solvent (H2O) = 0.05 kg

Hence,

Molality of solution = 0.8556 mol / 0.05 kg = 17.112 mol/kg = 17.1 m

Molality of solution = 17.1 m

 

Normality (N):

Normality is a measure of concentration equal to the gram equivalent weight per litre of solution. 

Normality = Mass of solute / (Volume of solution × Equivalent weight)

Equivalent weight = Molar mass / n

n = 1 (for NaCl)

Equivalent weight = 58.44 g mol-1 / 1 = 58.44 g mol-1

Hence,

Normality of solution = 50 g / (0.0633 L × 58.44 g mol-1) = 13.516 N = 13.5 N

Normality of solution = 13.5 N


Answers:

Mass % of solute = 50%

Molarity of solution = 13.5 M

Molality of solution = 17.1 m

Normality of solution = 13.5 N